OpenStudy (daffy_duck):

integrate cos(x)/(1+sin(x))....

OpenStudy (daffy_duck):



OpenStudy (anonymous):

Have you tried making the substitution \(u=1+\sin x\)?

OpenStudy (daffy_duck):

I don't know what the du would be.. My teacher never taught the trig integral stuff

OpenStudy (anonymous):

if u = 1 + sin(x) du = cos(x) dx plug that in and you get \[\int\limits_{}^{}\frac{ cosx }{1 + sinx }dx = \int\limits_{}^{}\frac{ du }{ u } = \ln|u| + C\] plug back in u = 1+ sinx \[= \ln|1+ \sin(x)| + C \]